OpenStudy (swaugh):

Which set of side lengths can form a right triangle? six, seven, eight ten, fourteen, thirty-six twelve, thirty-five, thirty-seven thirteen, twenty-four, thirty

OpenStudy (steve816):

Use the Pythagorean theorem.

OpenStudy (swaugh):

i dont know how to use Pythagorean theorem.

OpenStudy (steve816):

For example, the first one would be 6^2 + 7^2 = 8^2 And it doesn't match, so the first one is not right.

OpenStudy (steve816):

The Pythagorean theorem is simply \[a^2+b^2=c^2\]

OpenStudy (anonymous):

\[a ^{2}+b ^{2}=c^{2}\] c is the longest part of the triangle

OpenStudy (steve816):

ninja'd

OpenStudy (anonymous):

man steve816, you got faster then me

OpenStudy (swaugh):

so i add exponent 2 to all the numbers